SX LUT

18:47
  • TT mirror doesn't seem to be modulating.
  • Dataset 183354

20:30 Had communication problem to right box
  • Elliot unplug 110V from side of AGw1

21:00 Now TT mirror works (??)
  • just lucky?

21:13 IRTC darks
  • 20171221 dir
  • 14, 15

21:45 Start script
  • -90, 16, 17
  • -60, 18, 19
  • -30, 20, 21
  • 214632-214725 no good
  • 22-25 no good

21:50

22:00 Stop

-- %USERSIG{GregoryTaylor - 2017-12-21}%
Topic revision: r2 - 21 Dec 2017, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback